# Faster Model Predictive Control via Self-Supervised Initialization Learning

### v2

@@ IV-A Experiment Setup
The MPC objective function is composed of five parts as shown in Equation (7). The first two terms are $xte$ and Error in Heading ($eth$). $xte$ is computed using Equation 5. $eth$ is computed using Equation (8), which denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems. $v_{t}^{\text{ref}}$ is the desired speed of the vehicle. The last two terms regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother. $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term. The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds, which means that the MPC looks 0.5 seconds ahead. Since the calculation of $xte$ involves finding the closest point on a reference trajectory to the car’s current position, it introduces discontinuities and non-differentiability into the MPC objective function. As a result, only gradient-free, sampling-based solvers are suitable for this problem [27, 28]. Here we choose to use the gradient-free solver COBYLA [29] in our experiments.
Furthermore, given the high speed of the vehicle and the presence of multiple sharp turns on the track, gradient-free MPC solvers without a good initial guess of the control solution will struggle to optimize the trajectory in real time, making real-world implementation challenging. And we show in Section IV-C gradient-free MPC solver only succeeds in simple IMS track and fails at all of the complex tracks.
@@
We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$ to avoid MPC wasting time on optimizing a trajectory that is already good enough. Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.05 seconds in worst-case scenarios. As, 20Hz running rate of MPC is validated by [30] for vehicle path tracking.
(a) Training tracks
(b) Zero-shot tracks (complex)
(c) Zero-shot tracks (simple)
Figure 3: Tracks tested during training and testing.
IV-B Baselines and Metrics
@@
During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step. We compare our algorithm’s performance against five types of initial guess policies: All Zero: all-zero initial guesses, Previous Solution: initial guesses derived from the MPC solution at the previous step, Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training, Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC, and Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger. To ensure a fair comparison between the Ours w/o BC and warm-start policy trained by our two-phase learning algorithm, we extend the training time of Ours w/o BC by an additional hour to account for the data collection time in the offline BC phase.
IV-C Experiment Results
@@
We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)). The results are shown in Figure 4 and Figure 5, respectively. On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time and improves tracking accuracy. Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely through offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality covariance shift problem caused by BC.
[…]
@@
The results also reveal that initialize an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks. The vehicle consistently deviates from the lane when facing any sharp turns. This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track. Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively. This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
Figure 4: Experiment results on training tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
Figure 5: Experiment results on the six zero-shot testing tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
Furthermore, the Ours w/o RL algorithm also fails on both training and zero-shot domains. This is because the absence of offline BC requires the algorithm to explore the solution space from scratch, causing the online learning algorithm to take significantly longer to converge to a satisfactory solution. Given the same amount of training time, the quality of the policy learned by Ours w/o RL lags behind that of the policy learned through the two-phase training algorithm. This demonstrates the effectiveness of offline BC in accelerating online training.
To further inspect the relative performance of our approach when our baselines performed better, i.e. on simpler tracks, we conducted further testing on six warm-start policies on a simpler zero-shot track, IMS, as depicted in Figure 3(c),   similar to the path-tracking scenario in [31]. The results are summarized in Figure 6. Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
Figure 6: Experiment results on the simple IMS track.

### v3

@@ IV-A1 MPC Details
The MPC objective function is composed of five parts as shown in Eq. (5). The first two terms are Cross Track Error (xtexte) and Error in Heading (etheth) computed using Eq. (6). and Eq. (7) respectively. etheth denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems. $v_{t}^{\text{ref}}=10m/s$ is the desired speed of the vehicle. The last two terms regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother. $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term. The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds.
@@
Gradient-based solvers are highly effective when both system dynamics and cost functions are smooth and differentiable. However, in our setting, the calculation of xtexte (cross-track error) requires finding the closest point on a reference trajectory to the vehicle’s current position, which introduces discontinuities and non-differentiability into the objective function. This motivates the use of gradient-free solvers such as COBYLA [22], which are better suited for non-differentiable optimization problems [27, 28]. Nonetheless, with high speeds and on tracks with sharp turns, gradient-free MPC solvers struggle to optimize trajectories in real time without a good initial guess, making practical deployment challenging. As shown in Section V, the gradient-free MPC solver succeeds only on the simple IMS track but fails on all complex tracks.
We implement an early-stop condition when the planned trajectory’s accumulated xtexte is below $0.1,m$, preventing MPC wasting time on optimizing a trajectory that is already good enough. We also cap MPC iterations at 100 during RL training to avoid solver stagnation and evaluate performance under both 100 and 200 iteration limits during testing for sensitivity analysis.
IV-A2 RL Details
The reward at each step of the RL training is shown in Eq. (8). The first term is the negative MPC optimization time, and the second term is the negative of the xtexte over the planning horizon, $H$. This reward design helps optimize $\pi_{\theta}^{warm}$ by minimizing the MPC running time directly while maximizing tracking accuracy.
$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (8)
IV-B Sampling Based MPC
IV-B1 MPC Details
Unlike deterministic MPC, MPPI employs a sampling-based stochastic optimization framework to minimize a predefined cost. This setup allows us to test whether our warm-start algorithm can also enhance a different control optimizer. As MPPI is mainly used for path-planning and obstacle-avoidance tasks [29, 30], we build an obstacle-rich navigation domain for evaluation, as shown in Fig. 3. The training set includes two straight and two curved-road maps, while the testing set contains three unseen vertical and three unseen curved-road maps. Each map features randomly generated polygonal obstacles with varying sizes, shapes, and positions, posing diverse planning challenges. The vehicle starts at a green dot and must reach the goal region (red dot) without collisions. To increase training diversity, each training map offers two initialization points: one on the left and one on the right.
$J=\sum_{i=t}^{t+H-1}$ $\Big[w_{1}\,throttle_{i}^{2}+w_{2}\,steer_{i}^{2}$ (9)
$\quad+w_{3}\,(v_{i}-v_{t}^{ref})^{2}$
$\quad+w_{4}\,C_{obstacle}(x_{i},y_{i})$
$\quad+w_{5}\,C_{border}(x_{i},y_{i})\Big]$
$+\quad\sum_{i=t+H-3}^{t+H-1}\Big[w_{6}\,C_{goal}(x_{i},y_{i})\Big]$
The vehicle dynamics model is identical to that employed in the deterministic MPC in Eq. (4), ensuring a fair comparison. The cost function for MPPI is shown in Eq. (9). The first two terms correspond to smoothness penalties to enhance control stability. $C_{obstacle}$ penalizes proximity to obstacles, $C_{border}$ penalizes deviations beyond road boundaries to maintain lane centering, and $C_{goal}$ penalizes the final distance to the target region. The planning horizon is fixed at 25 steps. And number of samples is set to 200 samples to balance efficiency and control quality.
(a) Training Maps
(b) Testing Maps
Figure 3: Training and testing maps in the obstacle-rich navigation domain. Green dots indicate spawn points, red dots represent destinations, and polygons correspond to obstacles.
IV-B2 RL Details
The RL reward function is defined in Eq. (10). It is aligned with the MPPI objective function. The only difference is the final term, which encourages the vehicle to reduce its distance to the goal at each timestep. Since MPPI operates via parallel trajectory sampling rather than iterative optimization, we exclude the MPC optimization time from the RL reward, as all trajectories are sampled simultaneously on the GPU. However, warm-starting remains valuable for onboard systems with limited parallelism. A good initial guess allows the controller to achieve comparable performance with fewer samples or higher accuracy with the same number of samples. In our evaluation, we fix the number of samples and test whether our method outperforms the baselines in control performance.
$r_{t}^{mppi}$ $=-w_{1}\,throttle_{t}^{2}-w_{2}\,steer_{t}^{2}$ (10)
$\quad-w_{3}\,(v_{t}-v_{t}^{ref})^{2}-w_{4}\,C_{obstacle}(x_{t},y_{t})$
$\quad-w_{5}\,C_{border}(x_{t},y_{t})$
$\quad-w_{6}\,((C_{goal}(x_{t},y_{t})-(C_{goal}(x_{t-1},y_{t-1}))$
TABLE I: Deterministic MPC results with a maximum of 100 optimization iterations. Values represent averages over three test runs with different random seeds. Reported percentages indicate relative improvement compared to the corresponding baseline. The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s. means no significance.
All Zeros &
Previous Solution &
Ours w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Training Tracks
Zandvoort 0.1270 ± 0.0800 0.5929 ± 0.6066 0.1530 ± 0.0542 0.3869 ± 0.2840 0.1087 ± 0.0821 0.3975 ± 0.3084
Sochi 0.1429 ± 0.0999 0.5657 ± 0.5869 0.1772 ± 0.0620 0.4704 ± 0.5150 0.1128 ± 0.1010 0.3755 ± 0.3075
Hockenheim Failed 0.1129 ± 0.0912 0.4686 ± 0.4177 0.1450 ± 0.0581 0.4078 ± 0.4665 0.0865 ± 0.0877 0.3548 ± 0.3373
Training Tracks
Total
0.1285 ± 0.0919 0.5450 ± 0.5504 0.1597 ± 0.0600 0.4246 ± 0.4385 0.1033 ± 0.0920 0.3758 ± 0.3179
Compared with
Ours w/o RL
-24.27%
n.s.
22.10%
19.65%
30.04%
Zero-shot Tracks
Nuerburgring 0.1845 ± 0.0669 1.6630 ± 1.5488 0.1679 ± 0.0601 0.3774 ±0.3233 0.1158 ± 0.0967 0.4090 ± 0.3500
BrandsHatch 0.1163 ± 0.0922 0.5117 ± 0.5484 0.1545 ± 0.0519 0.3463 ± 0.2075 0.0895 ± 0.0837 0.3355 ± 0.2592
Oschersleben 0.1195 ± 0.0832 0.4181 ± 0.3514 0.1421 ± 0.0582 0.4383 ± 0.4309 0.1116 ± 0.0846 0.3405 ± 0.2395
Shanghai Failed 0.1845 ± 0.0732 1.3278 ± 1.2934 0.1717 ± 0.0602 0.4242 ± 0.4157 0.1189 ± 0.0977 0.3869 ± 0.2688
Melbourne 0.1375 ± 0.0963 0.7781 ± 0.9978 0.1767 ± 0.0588 0.4297 ± 0.3378 0.1235 ± 0.0937 0.3921 ± 0.3028
Catalunya 0.1140 ± 0.0942 0.5561 ± 0.6710 0.1561 ± 0.0571 0.4095 ± 0.3798 0.1019 ± 0.0891 0.3631 ± 0.2863
Zero-shot Tracks
Total
0.1558 ± 0.0871 1.0696 ± 1.2447 0.1629 ± 0.0589 0.4033 ± 0.3539 0.1095 ± 0.0913 0.3707 ± 0.2905
Compared with
Ours w/o RL
-4.57%
n.s.
62.30%
29.75%
65.34%
TABLE II: Deterministic MPC results with a maximum of 200 optimization iterations. Values represent averages over three test runs with different random seeds. Reported percentages indicate relative improvement compared to the corresponding baseline. The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s. means no significance.
All Zeros &
Previous Solution &
Ours w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Training Tracks
Zandvoort 0.2395 ± 0.1700 0.5130 ± 0.5054 0.2498 ± 0.1412 0.3472 ± 0.2664 0.1954 ± 0.1753 0.3501 ± 0.2402
Sochi 0.2821 ± 0.2079 0.5367 ± 0.5056 0.2836 ± 0.1649 0.3783 ± 0.3697 0.2166 ± 0.2122 0.3621 ± 0.2745
Hockenheim Failed 0.2072 ± 0.1914 0.4245 ± 0.3663 0.2349 ± 0.1502 0.3754 ± 0.4123 0.1593 ± 0.1815 0.3172 ± 0.2734
Training Tracks
Total
0.2451 ± 0.1935 0.4944 ± 0.4695 0.2578 ± 0.1545 0.3674 ± 0.3546 0.1918 ± 0.1930 0.3441 ± 0.2645
Compared with
Ours w/o RL
-5.19%
n.s.
25.70%
21.77%
30.41%
Zero-shot Tracks
Nuerburgring 0.3086 ± 0.1865 0.5814 ± 0.5536 0.2685 ± 0.1576 0.3359 ±0.3251 0.2358 ± 0.2008 0.4254 ± 0.3867
BrandsHatch 0.2172 ± 0.1887 0.4122 ± 0.3194 0.2514 ± 0.1421 0.3063 ± 0.1850 0.1584 ± 0.1771 0.3213 ± 0.2282
Oschersleben 0.2340 ± 0.1754 0.3977 ± 0.3196 0.2467 ± 0.1473 0.3623 ± 0.2838 0.2376 ± 0.1756 0.3893 ± 0.2840
Shanghai Failed 0.3093 ± 0.1943 0.6666 ± 0.6522 0.2806 ± 0.1551 0.3402 ± 0.3107 0.2150 ± 0.2028 0.3765 ± 0.2549
Melbourne 0.2747 ± 0.2011 0.6036 ± 0.6359 0.2930 ± 0.1561 0.3868 ± 0.3315 0.2223 ± 0.2016 0.3584 ± 0.2459
Catalunya 0.2300 ± 0.1947 0.5820 ± 0.6688 0.2663 ± 0.1521 0.3896 ± 0.3693 0.1920 ± 0.1871 0.3704 ± 0.3305
Zero-shot Tracks
Total
0.2663 ± 0.1949 0.5549 ± 0.5710 0.2697 ± 0.1531 0.3543 ± 0.3112 0.2078 ± 0.1921 0.3689 ± 0.2909
Compared with
Ours w/o RL
-1.30%
n.s.
36.16%
21.97%
33.52%
IV-C Baselines and Metrics
For deterministic MPC, we evaluate performance using two metrics: (1) the average optimization time per step (seconds) and (2) the average xtexte (meters) per step. For sampling-based MPC, we employ three metrics: (1) steering standard deviation, to assess control stability; (2) average number of steps required to reach the goal, as a proxy for path efficiency; and (3) number of obstacle collisions, to evaluate safety and avoidance capability. During testing, we compare the performance of our algorithm against five types of initial-guess policies:
  • All Zero: all-zero initial guesses.
  • Previous Solution: initial guesses derived from the MPC solution at the previous step.
  • Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training.
  • Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC.
  • Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger.
@@ IV-C Baselines and Metrics
To ensure a fair comparison between the Ours w/o BC and warm-start policy trained by our two-phase learning algorithm, we extend the training time of Ours w/o BC by an additional hour to account for the data collection time in the offline BC phase.
TABLE III: Sampling Based MPC Results. Values represent averages over three test runs with different random seeds. Numbers in parentheses denote the standard error.
All Zeros Previous Solution
Ours
w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Map
Name
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Training Maps
Vertical_1 0.6647
287.3333
± 22.8356
2.0000
± 1.4142
0.6306
251.8333
± 12.0568
0.0000
± 0.0000
0.4881
258.0000
± 19.7281
1.0000
± 1.0954
0.4747
278.1667
± 16.7978
0.6667
± 0.8165
0.4712
245.1667
± 20.8750
0.8333
± 1.6021
0.4689
249.3333
± 22.2860
0.0000
± 0.0000
Vertical_2 0.6491
289.0000
± 45.4093
0.6667
± 1.2111
0.6393
273.6667
± 11.5873
0.0000
± 0.0000
0.4923
264.0000
± 23.0304
0.8333
± 0.9832
0.4637
279.8333
± 24.7258
0.0000
± 0.0000
0.4810
265.6667
± 19.2942
0.0000
± 0.0000
0.4716
256.5000
± 26.8384
0.1667
± 0.4082
Curved_1 0.6290
369.8333
± 37.0751
1.5000
± 3.6742
0.6009
360.3333
± 13.4263
0.0000
± 0.0000
0.4773
376.3333
± 84.5167
0.0000
± 0.0000
0.4748
354.6667
± 36.9035
0.0000
± 0.0000
0.4322
334.6667
± 4.1793
0.0000
± 0.0000
0.4256
335.0000
± 5.6569
0.0000
± 0.0000
Curved_2 0.6366
422.6667
± 65.9019
9.3333
± 17.5803
0.6369
401.3333
± 26.1738
0.0000
± 0.0000
0.5064
427.6667
± 60.8857
1.0000
± 2.4495
0.4879
460.5000
± 81.4389
0.0000
± 0.0000
0.4812
375.6667
± 24.7117
1.3333
± 2.0656
0.4800
376.1667
± 17.6682
0.0000
± 0.0000
Training Map
Total
0.6433
342.2083
± 72.2255
3.3750
± 9.1358
0.6266
321.7917
± 64.5587
0.0000
± 0.0000
0.4922
331.5000
± 89.9633
0.7083
± 1.3981
0.4775
343.2917
± 87.7870
0.1667
± 0.4815
0.4663
305.2917
± 56.4088
0.5417
± 1.3507
0.4619
304.2500
± 57.5917
0.0417
± 0.2041
Testing Maps
Vertical_1 0.6425
285.6667
± 107.7698
9.6667
± 7.7675
0.6228
288.3333
± 97.6183
0.0000
± 0.0000
0.4667
225.0000
± 1.0000
0.0000
± 0.0000
0.4982
355.6667
± 198.4574
0.0000
± 0.0000
0.4697
230.0000
± 21.7945
0.3333
± 0.5774
0.4579
220.6667
± 4.5093
0.0000
± 0.0000
Vertical_2 0.6416
266.3333
± 35.5575
2.0000
± 2.6458
0.6373
231.6667
± 3.0551
0.0000
± 0.0000
0.5068
235.6667
± 14.2244
1.0000
± 1.0000
0.4709
225.3333
± 4.1633
0.6667
± 1.1547
0.4592
217.6667
± 3.2146
0.0000
± 0.0000
0.4421
219.3333
± 14.4338
0.0000
± 0.0000
Vertical_3 0.6298
312.0000
± 106.2262
3.3333
± 4.9329
0.6280
263.6667
± 38.5011
0.0000
± 0.0000
0.4332
220.3333
± 1.1547
0.0000
± 0.0000
0.4646
258.0000
± 4.3589
0.0000
± 0.0000
0.4685
257.3333
± 1.5275
0.0000
± 0.0000
0.4621
239.0000
± 25.1197
0.0000
± 0.0000
Curved_1 0.6251
348.0000
± 2.6458
9.6667
± 12.6623
0.5929
328.6667
± 3.0551
0.0000
± 0.0000
0.4457
323.6667
± 3.5119
0.0000
± 0.0000
0.4385
328.0000
± 10.5357
0.0000
± 0.0000
0.4267
317.6667
± 0.5774
0.0000
± 0.0000
0.4227
315.6667
± 1.1547
0.0000
± 0.0000
Curved_2 0.5890
331.0000
± 9.1652
0.3333
± 0.5774
0.5720
319.6667
± 0.5774
0.0000
± 0.0000
0.4163
314.0000
± 1.7321
0.0000
± 0.0000
0.4060
310.6667
± 1.1547
0.0000
± 0.0000
0.4130
308.6667
± 2.0817
0.0000
± 0.0000
0.4120
310.0000
± 4.5826
0.0000
± 0.0000
Curved_3 0.6416
418.0000
± 83.5165
3.3333
± 3.5119
0.5841
339.3333
± 8.1445
0.0000
± 0.0000
0.4654
340.6667
± 26.3881
0.0000
± 0.0000
0.4975
387.6667
± 69.8164
1.3333
± 2.3094
0.4446
324.3333
± 5.6862
0.0000
± 0.0000
0.4330
326.0000
± 7.2111
0.0000
± 0.0000
Testing Map
Total
0.6284
326.8333
± 78.7821
4.7222
± 6.7196
0.6050
295.2222
± 53.3728
0.0000
± 0.0000
0.4550
276.5556
± 52.8597
0.1667
± 0.5145
0.4713
310.8889
± 91.9545
0.3333
± 1.0290
0.4467
275.9444
± 44.7667
0.0556
± 0.2357
0.4372
271.7778
± 48.6047
0.0000
± 0.0000
TABLE IV: Relative improvement between different methods in sampling-based MPC and statistical test results. The reported percentages indicate relative improvement. The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s. means no significance.
Previous Solution
Ours
w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Training Maps
Compared with
All Zeros
2.60%
5.97%
100%
23.49%
3.13%
n.s.
79.01%
25.77%
-0.32%
n.s.
95.06%
27.51%
10.79%
83.95%
28.20%
11.09%
98.76%
Compared with
Previous Solution
21.45%
-3.02%
n.s.
23.80%
-6.68%
n.s.
25.58%
5.13%
26.27%
5.46%
Compared with
Ours w/o BC
2.99%
-3.56%
n.s.
76.46%
5.26%
7.91%
23.52%
n.s.
6.16%
8.22%
94.11%
Compared with
Ours w/o RL
2.35%
n.s.
11.07%
-224.96%
n.s.
3.27%
n.s.
11.37%
74.99%
n.s.
Compared with
Ours w/o DAgger
0.95%
n.s.
0.34%
n.s.
92.30%
Testing Maps
Compared with
All Zeros
3.72%
9.67%
100.00%
27.59%
15.38%
96.47%
25.00%
4.88%
92.94%
28.91%
15.57%
98.82%
30.43%
16.85%
100.00%
Compared with
Previous Solution
24.79%
6.32%
22.10%
-5.31%
n.s.
26.17%
6.53%
27.74%
7.94%
0.00%
Compared with
Ours w/o BC
-3.58%
n.s.
-12.41%
n.s.
-99.94%
n.s.
1.82%
n.s.
0.22%
n.s.
66.65%
n.s.
3.91%
1.73%
100.00%
n.s.
Compared with
Ours w/o RL
5.21%
n.s.
11.24%
83.32%
n.s.
7.22%
12.58%
100.00%
n.s.
Compared with
Ours w/o DAgger
2.13%
1.51%
n.s.
100.00%
n.s.
V Experiment Results
In this section we design experiments to test the following research questions in both deterministic MPC and sampling-based MPC problem:
  • RQ1 Does our warm-start policy reduce solver runtime in deterministic MPC problem?
  • RQ2 Does our warm-start policy improve tracking performance in deterministic MPC problem?
  • RQ3 Does our warm-start policy generalize to unseen tracks in deterministic MPC problem?
  • RQ4 Does our warm-start policy generalize from deterministic MPC to sampling-based MPC problem?
V-A Deterministic MPC
@@
We perform testing on both the training tracks (Fig. 2(a)) and the challenging zero-shot tracks (Fig. 2(b)). The results are shown in Table I. On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time and improves tracking accuracy. Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely via offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality and covariance shift problem in BC.
On the training tracks, compared to Ours w/o RL, the proposed method improves optimization time by $19.95\%$ and tracking accuracy by $30.60\%$. When compared to Ours w/o RL, Ours w/o DAgger achieves a $19.12\%$ improvement in tracking accuracy; however, the optimization time degrades by $22.08\%$, highlighting the necessity of incorporating the DAgger training loss during RL training.
On the zero-shot tracks, compared to Ours w/o RL, the proposed method improves optimization time by $21.63\%$ and tracking accuracy by $34.12\%$. Similarly, Ours w/o DAgger achieves a $30.35\%$ improvement in tracking accuracy, but optimization time degrades by $18.00\%$. This degradation pattern aligns with that on the training tracks, further highlighting the importance of integrating the DAgger training loss during RL training.
Figure 4: Visualization of optimized trajectories for our method and the baselines on a segment of the Nürburgring map.
@@
The results also reveal that initializing an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks. The vehicle consistently deviates from the lane when facing any sharp turns. This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track. Since the real-time MPC only optimizes the solution for a maximum of 100 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively. This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
We also relaxed the maximum MPC iterations to 200 and repeated the test. The results in Table II show that our warm-start policy again achieves the greatest improvement in optimization time and the largest performance gains over the baselines, further supporting RQ1, RQ2, and RQ3.
A visualization of the optimized trajectories for our method and the baselines on a segment of the Nürburgring map is presented in Fig. 4 (MPC optimization iterations=100). The figure illustrates that using our method as a warm-start strategy enables the MPC to track the reference path more accurately by tracking the sharp more accurately.
V-B Sampling-Based MPC
We evaluate performance on the training maps (Fig. 3(a)) and the zero-shot maps (Fig. 3(b)), with results summarized in Table III. Consistent with the findings for deterministic MPC, our proposed framework delivers the most significant performance improvements for the MPPI controller compared to all baselines. Notably, the improvements are more pronounced in the zero-shot environments, demonstrating the strong generalizability of the proposed method.
V-C Findings